Adam is 5 years younger than Eve. In 1 year, Even will be three times as old as Adam was 4years ago. Find there ages now.
Answer by checkley71(8403)

EVE=X
ADAM=(X-5) AGES NOW.
IN 1 YEAR EVE WILL BE (X+1)
IN 1 YEAR ADAM WILL BE (X-5)+1
(X+1)=3(X-9)
X+1=3X-27
3X-X=27+1
2X=28
X=14 AGE OF EVE NOW.
14-5=9 AGE OF ADAM NOW.
IN 1 YEAR EVE WILL BE 15
ADAM 4 YEARS AGO WAS 9-4=5
P[ROOF:
15=3(14-9)
15=3*5
15=15